# Adapting Vision-Language Foundation Model for Next Generation Medical Ultrasound Image Analysis

### v1

@@ 1 Introduction
To address this problem, this study proposes a simple and effective workflow for the foundation model domain adaptation from natural image distribution to ultrasound modality. Specifically, we optimized the pre-trained CLIP [28] in a parameter-efficient-fine-tuning (PEFT) manner by introducing a hierarchical adapter to the vision branch. To further enhance the visual representation capability for ultrasound images with ROIs in various characteristics, and reduce the complexity of whole architecture, a straightforward up-sampling and adaptive pooling layer are integrated to the foundation model as head for specific downstream tasks. The proposed framework is evaluated on six ultrasound datasets: the in-house lymph node ultrasound datasets from two different centers (LN-1, LN-2), the public available breast ultrasound image (BUSI) dataset [3], thyroid datasets DDTI [25] and TN3K [9], and ultrasound scans of prostate [15]. The experimental findings demonstrate that the proposed framework surpasses existing methods including the original version of CLIP and other variants which were trained from scratch based on medical datasets.
@@ 2 Related Work
Numerous studies have investigated the applications of FMs in medical image analysis. Huix et al. [14] evaluated FMs on medical image classification tasks and found that, while the these models can outperform state-of-the-art approaches pre-trained on ImageNet [6], their performance varies inconsistently across datasets of different scales, suggesting limitations in robustness and generalizability.
In contrast to leveraging original pre-trained weights, some researchers have pursued training CLIP-like models from scratch using exclusively medical data. PMC-CLIP [20] harnessed approximately one million biomedical image-text pairs from PubMed Central Open Access Subset across 12 modalities, achieving state-of-the-art results in various downstream tasks such as image-text retrieval, image classification and medical VQA. BioMedCLIP [38] further scaled this effort by expanding the closed-source PMC-15M dataset to 15 million pairs. To address the gap of data transparency, UniMedCLIP [17] curated an open-source cross-modal dataset to pre-train a unified vision-language model, which comprising over 5.3 million samples and across six modalities.
@@ 4.1 Datasets
There are three types of usage of datasets in this study, including fine-tuning (PMC-OA, CURD), segmentation and classification. The overview of datasets and partitions are shown in Table 1.
@@ 4.3 Perquisites
The state-of-the-art CLIP variants used for comparison in this study include BioMedCLIP [38], UniMedCLIP [17], and MetaCLIP [34]. To establish a visual-only performance baseline, we also evaluate models that rely solely on image input, including UNet [30], ResNet18 [12], and DINOv2 [24]. For all experiments and comparisons, we adopt the ViT-Base architecture with a patch size of 16$\times$16 pixels as the standard backbone. Pre-trained weights are used when publicly available for the respective models.
@@ 4.5 Classification Results
CLIP (LoRA, ft.) yields mixed but promising results and has made a notable improvement compares to the version that without fine-tuning process, gaining up to 3.64% increase across three datasets. These results suggest that LoRA adaptation combined with supervised fine-tuning can harness the pre-trained features of CLIP to approach state-of-the-art performance. This is particularly evident on datasets with more homogeneous lesion characteristics, while it may struggle when sample diversity is higher or data scarcity is more pronounced (as in LN-2).

### v2

@@ 1 Introduction
To address this problem, this study proposes a simple and effective workflow for the foundation model domain adaptation from natural image distribution to ultrasound modality. Specifically, we optimized the pre-trained CLIP [28] in a parameter-efficient-fine-tuning manner by introducing a hierarchical adapter to the vision branch. To further enhance the visual representation capability for ultrasound images with ROIs in various characteristics, and reduce the complexity of whole architecture, a straightforward up-sampling and adaptive pooling layer are integrated to the foundation model as head for specific downstream tasks. The proposed framework is evaluated on six ultrasound datasets: the in-house lymph node ultrasound datasets from two different centers (LN-1, LN-2), the public available breast ultrasound image (BUSI) dataset [3], thyroid datasets DDTI [25] and TN3K [9], and ultrasound scans of prostate [15]. The experimental findings demonstrate that the proposed framework surpasses existing methods including the original version of CLIP and other variants which were trained from scratch based on medical datasets.
@@ 2 Related Work
Numerous studies have investigated the applications of FMs in medical image analysis. Huix et al. [14] evaluated FMs on medical image classification tasks and found that, while the these models can outperform state-of-the-art (SOTA) approaches pre-trained on ImageNet [6], their performance varies inconsistently across datasets of different scales, suggesting limitations in robustness and generalizability.
In contrast to leveraging original pre-trained weights, some researchers have pursued training CLIP-like models from scratch using exclusively medical data. PMC-CLIP [20] harnessed approximately one million biomedical image-text pairs from PubMed Central Open Access Subset across 12 modalities, achieving SOTA results in various downstream tasks such as image-text retrieval, image classification and medical VQA. BioMedCLIP [38] further scaled this effort by expanding the closed-source PMC-15M dataset to 15 million pairs. To address the gap of data transparency, UniMedCLIP [17] curated an open-source cross-modal dataset to pre-train a unified vision-language model, which comprising over 5.3 million samples and across six modalities.
@@ 4.1 Datasets
There are three types of usage of datasets in this study, including fine-tuning (PMC-OA [20], CURD [19]), segmentation and classification. The overview of datasets and partitions are shown in Table 1.
@@ 4.3 Perquisites
The SOTA CLIP variants used for comparison in this study include BioMedCLIP [38], UniMedCLIP [17], and MetaCLIP [34]. To establish a visual-only performance baseline, we also evaluate models that rely solely on image input, including UNet [30], ResNet18 [12], and DINOv2 [24]. For all experiments and comparisons, we adopt the ViT-Base architecture with a patch size of 16$\times$16 pixels as the standard backbone. Pre-trained weights are used when publicly available for the respective models.
@@ 4.5 Classification Results
CLIP (LoRA, ft.) yields mixed but promising results and has made a notable improvement compares to the version that without fine-tuning process, gaining up to 3.64% increase across three datasets. These results suggest that LoRA adaptation combined with supervised fine-tuning can harness the pre-trained features of CLIP to approach SOTA performance. This is particularly evident on datasets with more homogeneous lesion characteristics, while it may struggle when sample diversity is higher or data scarcity is more pronounced (as in LN-2).